Here are some historical data on the risk characteristics of Bank of America and Starbucks.

Bank of America Starbucks
β (beta) 1.68 .73
Yearly standard deviation of return (%) 30.7 17.0

Assume the standard deviation of the return on the market was 19%. (Use decimals, not percents, in your calculations.)

     

a. The correlation coefficient of Bank of America's return versus Starbucks is .44. What is the standard deviation of a portfolio invested half in each stock? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

Standard deviation             %

b. What is the standard deviation of a portfolio invested one-third in Bank of America, one-third in Starbucks, and one-third in risk-free Treasury bills? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

Standard deviation             %

c. What is the standard deviation if the portfolio is split evenly between Bank of America and Starbucks and is financed at 50% margin, that is, the investor puts up only 50% of the total amount and borrows the balance from the broker? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

Standard deviation             %

d-1. What is the approximate standard deviation of a portfolio comprised of 100 stocks with betas of 1.68 like Bank of America? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

Standard deviation             %

d-2. What is the approximate standard deviation of a portfolio comprised of 100 stocks with betas of .73 like Starbucks? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

Standard deviation             %
Answer #1

Standard deviation for portfolio of two stocks is calculated by = sqrt{(w1^{2} sigma 1^{2})+(w2^{2} sigma 2^{2})+ (2* w1* w2*sigma 1 *sigma 2* ho )}

where w1 and w2 are weights of stocks in the portfolio , σ1, σ2 are standard deviations of the stocks and ho is correlation between the two

a) sigma p = sqrt{(w1^{2}sigma 1^{2})+(w2^{2}sigma 2^{2})+ (2* w1* w2*sigma 1 *sigma 2* ho )}

= (0.5 +30.72)+ (0.52 +172)+(2 0.5+ 0.5+30.7*17*0.44) =20.5594 %

b) As treasury bills are risk free , their standard deviation = 0 . Also correlation between treasury bill and stocks = 0

Standard deviation for portfolio of three assets is calculated by =

sqrt{(w1^{2} sigma 1^{2})+(w2^{2} sigma 2^{2})+(w3^{2} sigma 3^{2})+ (2* w1* w2*sigma 1 *sigma 2* ho(1,2)) +(2* w2* w3*sigma 2*sigma 3* ho (2,3))+(2* w1* w3*sigma 1 *sigma 3* ho (1,3))}= (0.332 30.72)(0.332 17)+(0.332 *02)(20.33 0.33 30.7 17*0.44)(20.330.3330.7*00)(2 0.33 0.33 *0 17 0) =13.569%

c) Where 50% of the amount is financed , weight of risk free treasury bills = -50%

so weight of each stock can be given by equation w1+ w2 + (-50%) = 100% here w1 = w2 as the portfolio is split evenly between Bank of America and Starbucks.

2w =150% , w = 75%

Standard deviation for portfolio of three assets is calculated by =

sqrt{(w1^{2} sigma 1^{2})+(w2^{2} sigma 2^{2})+(w3^{2} sigma 3^{2})+ (2* w1* w2*sigma 1 *sigma 2* ho(1,2)) +(2* w2* w3*sigma 2*sigma 3* ho (2,3))+(2* w1* w3*sigma 1 *sigma 3* ho (1,3))}= V(0.75? * 30.72) + (0.75? * 172) + +(一0.5? * 02) (2 * 0.75 * 0.75 * 30.7 * 17 * 0.44) + (2 * 0.75 *一0.5 * 30.7 * 0 * 0) (2 * =30.839%

d) Standard deviation for portfolio of n similar assets = n * V ( (1/n)2 * σ2) = n* 1/n * sigma = sigma

d-1) standard deviation of a portfolio comprised of 100 stocks with betas of 1.68 like Bank of America = standard deviation of Bank of America = 30.7%

d-2) standard deviation of a portfolio comprised of 100 stocks with betas of .73 like Starbucks = standard deviation of Starbucks = 17%